OpenStudy (anonymous):

The radius of a circular puddle of water is increasing at a rate of 2.5 cm/s. Find the exact rate at which the area is increasing at the instant the radius is 12 cm.

OpenStudy (anonymous):

a = pi r^2 find da/dt you know that dr/dt = 2.5cm/s find da/dt for r = 12cm

OpenStudy (anonymous):

got it jennhy?

OpenStudy (anonymous):

let me know if you have doubts

OpenStudy (anonymous):

can you please check my solution?

OpenStudy (anonymous):

can you rewrite it, you seem to have used x and r for the same things, use r for radius and a for area.

OpenStudy (anonymous):

wouldnt that give me the same answer?

OpenStudy (anonymous):

\[\frac{da}{dt}=\frac{da}{dr} \times \frac{dr}{dt} = \frac{d \pi r^2}{dr} \times 2.5\]

OpenStudy (anonymous):

\[= 2 \pi r \times 2.5\]

OpenStudy (anonymous):

at r = 12, da/dt = 60 pi

OpenStudy (anonymous):

Sorry but im unsure of how you got dπr2dr×2.5

OpenStudy (anonymous):

using chain rule, or is it product rule? can't remember which is which, but I used that

OpenStudy (anonymous):

its the chain rule, what i da/dr?

OpenStudy (anonymous):

a = pi r^2 so da/ dr = d(pi r^2)/dr = 2 pi r

OpenStudy (anonymous):

OHHHHHH I GET IT NOW! THANK YOU :D

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

Feel free to solve my other questions if you dont mind hehe

OpenStudy (anonymous):

i wont solve them for you, but I can show you the method to follow. post them

OpenStudy (anonymous):

YES THEATS FINE :)

OpenStudy (anonymous):

*THATS

OpenStudy (anonymous):

A bushwalker can walk at 5 km/h through clear land and 3 km/h through bushland. If she has to get from point A to point B following a route indicated at right, find the value of x so that the route is covered in a minimum time. (Note: time = distance ) speed

OpenStudy (anonymous):